OpenStudy (anonymous):

Solve the following system of equations. Variables x,y,z. 7x-y+z=-22 3x+2y-3z=-27 x-3y+2z=11

OpenStudy (texaschic101):

7x - y + z = -22 --->(2) 3x + 2y - 3z = -27 ---------------- 14x - 2y + 2z = -44 (result of multiplying by 2) 3x + 2y - 3z = -27 -----------------add 17x - z = -71 3x + 2y - 3z = -27 --->(3) x - 3y + 2z = 11--->(2) -------------------- 9x + 6y - 9z = - 81 (result of multiplying by 3) 2x - 6y + 4z = 22 (result of multiplying by 2) -------------------add 11x - 5z = - 59 17x - z = -71 --->(-5) 11x - 5z = -59 ---------------- -85x + 5z = 355 (result of multiplying by -5) 11x - 5z = -59 ---------------add -74x = 296 x = - 296/74 x = - 4 17x - z = -71 17(-4) - z = -71 -68 - z = -71 -z = -71 + 68 -z = -3 z = 3 x - 3y + 2z = 11 -4 - 3y + 2(3) = 11 -4 - 3y + 6 = 11 -3y + 2 = 11 -3y = 11 - 2 -3y = 9 y = -3 check... 7x - y + z = -22 7(-4) - (-3) + 3 = -22 -28 + 3 + 3 = -22 -28 + 6 = -22 -22 = -22 (correct) x = -4, y = -3, and z = 3